# Thought Anchors: Which LLM Reasoning Steps Matter?

### v3

@@ Abstract
Reasoning large language models have recently achieved state-of-the-art performance in many fields. However, their long-form chain-of-thought reasoning creates interpretability challenges as each generated token depends on all previous ones, making the computation harder to decompose. We argue that analyzing reasoning traces at the sentence level is a promising approach to understanding reasoning processes. We present three complementary attribution methods: (1) a black-box method measuring each sentence’s counterfactual importance by comparing final answers across 100 rollouts conditioned on the model generating that sentence or one with a different meaning; (2) a white-box method of aggregating attention patterns between pairs of sentences, which identified “broadcasting” sentences that receive disproportionate attention from all future sentences via “receiver” attention heads; (3) a causal attribution method measuring logical connections between sentences by suppressing attention toward one sentence and measuring the effect on each future sentence’s tokens. Each method provides evidence for the existence of thought anchors, reasoning steps that have outsized importance and that disproportionately influence the subsequent reasoning process. These thought anchors are typically planning or backtracking sentences. We provide an open-source tool (thought-anchors.com) for visualizing the outputs of our methods, and present a case study showing converging patterns across methods that map how a model performs multi-step reasoning. The consistency across methods demonstrates the potential of sentence-level analysis for a deeper understanding of reasoning models.
@@ 1 Introduction
Figure 1: Summary of our three methods for principled attribution to important sentences in reasoning traces. A. An example reasoning trace with sentences labeled per our taxonomy. B. Our proposed methods are: black-box resampling, receiver heads, and attention suppression. C. A directed acyclic graph among sentences prepared by one of our techniques, made available open source.
@@
Training large language models to reason with chain-of-thought (Reynolds and McDonell, 2021; Nye et al., 2021; Wei et al., 2023) has led to significant advances in capabilities (OpenAI, 2024). The resulting reasoning traces are already used in safety research (Baker et al., 2025; Shah et al., 2025). However, there has been little work adapting interpretability methods to this new paradigm (though see (Venhoff et al., 2025; Goodfire, 2025)). Traditional mechanistic interpretability (Olah et al., 2020; Olah, 2022) methods often focus on a single forward pass of the model: understanding the activations, how they are processed by each layer, and how they are converted to the final output (Wang et al., 2022; Heimersheim and Janiak, 2023; Hanna et al., 2023). However, this framework is too fine-grained for autoregressive reasoning models, which consume their own output tokens.
@@
A core step in many interpretability strategies is to decompose the model into smaller parts that can be analyzed independently (Lindsey et al., 2025). A natural decomposition for chain-of-thought is into individual sentences and how they depend on each other. Interpretations of neural network behavior operate at varying levels of abstraction (Geiger et al., 2021, 2025), and sentence-level explanations strike an intermediate abstraction depth. Compared to tokens, sentences are more coherent and often coincide with reasoning steps extracted by an LLM (Venhoff et al., 2025; Arcuschin et al., 2025). Compared to paragraphs, sentences are less likely to conflate reasoning steps and may serve as an effective target for linking different steps.
@@
Prior work has established that different sentences within reasoning traces perform distinct functions. Backtracking sentences (e.g., “Wait…”) cause the model to revisit earlier conclusions, which boosts final-answer accuracy (Muennighoff et al., 2025). Other research has distinguished sentences based on whether they retrieve new information or execute deduction with existing information (Venhoff et al., 2025). Hence, reasoning may follow an overarching structure, where sentences can introduce and pursue high-level computational goals. Yet, approaches for mapping this structure remain limited.
We propose three complementary methods for mapping the structure of reasoning traces that focus on what we term thought anchors: critical reasoning steps that guide the rest of the reasoning trace. We provide evidence for this type of anchoring based on black-box evidence from resampling and white-box evidence based on attention patterns.
@@
First, in section˜3 we present a black-box method for measuring the counterfactual impact of a sentence on the model’s final answer and future sentences. We repeatedly resample reasoning traces from the start of each sentence. Based on resampling data, we can quantify the impact of each sentence on the likelihood of any final answer or the likelihood of producing any subsequent sentence. Via this resampling approach, we can additionally distinguish planning sentences that initiate computations leading to some answer from sentences performing computations necessary for the answer but which are predetermined.
Second, in section˜4 we present a white-box method for evaluating importance based on the sentences most attended. Our analyses reveal “receiver” heads that narrow attention toward particular past “broadcasting” sentences. Compared to base models, where attention is more diffuse, reasoning models display overall greater attentional narrowing through receiver heads, and these heads have an outsized impact on the model’s final answer. We develop a systematic approach to identifying receiver heads and show how evaluating sentences on the extent to which they are broadcast by these heads provides a mechanistic measure of importance.
@@
Third and finally, in section˜5 we present a method that measures the causal dependency between specific pairs of sentences in a reasoning trace. For each sentence in a trace, we intervene by masking all attention to it from subsequent tokens. We then measure the effect of suppression on subsequent token logits (KL divergence) compared to those generated during the absence of suppression. Averaging token effects by sentence, this strategy measures each sentence’s direct causal effect on each subsequent sentence.
@@
2 Setup
2.1 Sentence taxonomy for reasoning functions
We categorized sentences using the framework by Venhoff et al. (2025), which describes distinct reasoning functions within a reasoning trace. We define eight categories (see examples and frequencies in Appendix˜A):
@@ 2.1 Sentence taxonomy for reasoning functions
Each sentence in the analyzed response is assigned to one of these categories using an LLM-based auto-labeling approach. The prompt used to label sentence categories is provided in Appendix˜B. Some categories that rarely appear are omitted from the figures below. Probing classifiers achieve high accuracy in distinguishing sentences and their corresponding categories (see Appendix˜C).
2.2 Model and dataset
Our analyses are based on the DeepSeek R1-Distill Qwen-14B model (48 layers) (DeepSeek, 2025). We used a temperature of $0.6$ and a top-p value of $0.95$ in all of our experiments. We used the MATH dataset (Hendrycks et al., 2021) specifically targeting 10 challenging mathematics questions that the model correctly solves 25-75% of the time, identified by testing the model on 1,000 problems 10 times each. For each selected problem, we generated one correct and one incorrect reasoning trace, producing 20 responses. The average response is 144.2 sentences long (95% CI: [116.7, 171.8] and 4208 tokens (95% CI: [3479, 4937]). We focus only on sentences before the resampling method indicates the model has converged on an answer (i.e., after which it gives the same response in >98% of resamples). In Appendix˜D, we provide results from applying our techniques to other reasoning models.
3 Measuring counterfactual influence
3.1 Forced answer importance
[…]
The taxonomic categories reveal that active computation statements produce the greatest distributional shifts in this forced answer importance (Figure 3A). This aligns with prior findings that reasoning models update their beliefs throughout the trace, and suggests intermediate calculations alter the model’s answer responses. However, active computation steps may not provide a comprehensive view on importance, as they may be predetermined by upstream statements.
@@ 3.2 Counterfactual importance
A natural limitation of the forced-answer approach is that a sentence $S$ may be necessary for some final answer, but consistently produced by the LLM late in the reasoning trace. This means that forced answer accuracy will be low for all sentences that occur before $S$, not allowing the importance of these earlier steps to be accurately determined.
Consider a rollout which consists of sentences $S_{1},S_{2},\dots,S_{i},\dots,S_{M}$ and a final answer $A$. We use resampling to define a metric capturing the extent to which sentence $S$ causes $A$ to be incorrect. [^1] All $S_{j}$ and $T_{k}$ as well as $A$ in this section are random variables, since we sample with non-zero temperature. We call this metric counterfactual importance. We can motivate and define this metric with three steps:
@@
    Rollout sampling. For a given sentence $S_{i}$, we generate a distribution over final answers by generating 100 rollouts both without sentence $S_{i}$ (the base condition, with rollouts of the form $S_{1},S_{2},\dots,S_{i-1},T_{i},\dots,T_{N},A^{\prime}_{S_{i}}$), and another distribution with sentence $S$ (the intervention condition, with rollouts of the form $S_{1},S_{2},\dots,S_{i-1},S_{i},\dots,S_{M},A_{S_{i}}$).
@@
    Naive distributional comparison. We can compute the KL Divergence between the final answer distributions in the two conditions, i.e., $\text{importance}_{\text{r}}:=D_{\text{KL}}[p(A^{\prime}_{S_{i}})||p(A_{S_{i}})]$, providing a scalar measure of how much sentence $S_{i}$ changes the answer. We call this the resampling importance metric.
  3. 3.
@@
    Semantic filtering. The problem with resampling importance is that in rollout resampling, if $T_{i}$ is identical or similar to $S_{i}$ then we do not get much information about whether $S_{i}$ is important or not. Therefore, we write $S\not\approx T$ if two sentences $S$ and $T$ are dissimilar, defined as having cosine similarity less than the median value across all sentence pairs in our dataset, when fed into an embeddings model (see Appendix˜E for details). Therefore we can finally define the counterfactual importance by conditioning on $T_{i}\not\approx S_{i}$ i.e., $\text{importance}:=D_{\text{KL}}[p(A^{\prime}_{S_{i}}|T_{i}\not\approx S_{i})||p(A_{S_{i}})]$.
Figure 2: Accuracy over 100 rollouts at each sentence for (A) one correct and (B) one incorrect base solution. Red dots mark significant spikes or dips. Local minima and maxima sentences are annotated with category initials (e.g., PG = plan generation). The analyses below focus on the counterfactual KL-divergence between sentences, but resampling accuracy is visualized here as it is more intuitive.
@@
Since we resample all steps after a given sentence $S_{i}$, we avoid the aforementioned limitation of forced-answering. We also provide empirical evidence that the principled counterfactual importance definition in 1-3 above is useful, by comparing it to the resampling importance in Appendix˜G.
@@
Across our dataset, we find that plan generation and uncertainty management (e.g., backtracking) sentences consistently show higher counterfactual importance than other categories like fact retrieval or active computation (see Figure 3B). This supports the view that high-level organizational sentences anchor, organize, and steer the reasoning trajectory. We argue that this approach provides more informative results compared to forced answer importance and prior token- or attention-based metrics.
@@
Figure 3: The mean of each sentence category for (A) forced-answer importance and (B) counterfactual importance, per the resampling method, plotted against the sentence category’s mean position in the reasoning trace. Only the 5 most common sentence types are shown (see Appendix˜F).
3.3 Sentence-to-sentence importance
In addition to estimating a sentence’s effect on the final answer, we extend our framework to quantify its causal influence on each individual future sentence. This sentence-level attribution helps uncover the structure of the model’s chain-of-thought: which sentences initiate subroutines, influence future claims, or act as attractors for subsequent steps. We estimate the counterfactual importance of sentence $S_{i}$ on a future sentence $S^{\text{Fut.}}$ formally with:
In this context $S^{\text{Fut.}}$ semantically occurs if, when we extract the sentences and identify the best candidate match for $S^{\text{Fut.}}$ using cosine similarity between sentence embeddings, it has greater than $0.8$ cosine similarity (i.e., the median value in our dataset) to that sentence. Pseudocode for estimating sentence-to-sentence importance and empirical values of this metric can be found in Appendix˜G.
4 Attention aggregation
@@
We hypothesize that important sentences may receive heightened attention from downstream sentences. Although attention weights do not necessarily imply causal links (see Section 5), heightened attention is a plausible mechanism by which important sentence may exert their influence on subsequent sentences. We conjecture further that a high focus on important sentences may be driven by specific attention heads, and by tracking such heads, we may pinpoint key sentences.
@@
We assessed the degree different heads narrow attention toward particular sentences. First, for each reasoning trace, we averaged each attention head’s token-token attention weight matrix to form a sentence-sentence matrix, where each element is the mean across all pairs of tokens between two sentences. Based on each attention matrix, we computed the mean of its column below the diagonal to measure the extent each sentence receives attention from all downstream sentences; averaged only among pairs at least four sentences apart to focus on distant connections. This generates a distribution for each head (e.g., Figure 4A), and the extent each head generally narrows attention toward specific sentences can be quantified as its distribution’s kurtosis (computed for each reasoning trace, then averaged across traces). Plotting each head’s kurtosis reveals that some attention heads strongly narrow attention toward specific, possibly important, sentences in the reasoning trace (Figure 4B).

### v4

@@ Abstract
Current frontier large-language models rely on reasoning to achieve state-of-the-art performance. Many existing interpretability are limited in this area, as standard methods have been designed to study single forward passes of a model rather than the multi-token computational steps that unfold during reasoning. We argue that analyzing reasoning traces at the sentence level is a promising approach to understanding reasoning processes. We introduce a black-box method that measures each sentence’s counterfactual importance by repeatedly sampling replacement sentences from the model, filtering for semantically different ones, and continuing the chain of thought from that point onwards to quantify the sentence’s impact on the distribution of final answers. We discover that certain sentences can have an outsized impact on the trajectory of the reasoning trace and final answer. We term these sentences thought anchors. These are generally planning or uncertainty management sentences, and specialized attention heads consistently attend from subsequent sentences to thought anchors. We further show that examining sentence-sentence causal links within a reasoning trace gives insight into a model’s behavior. Such information can be used to predict a problem’s difficulty and the extent different question domains involve sequential or diffuse reasoning. As a proof-of-concept, we demonstrate that our techniques together provide a practical toolkit for analyzing reasoning models by conducting a detailed case study of how the model solves a difficult math problem, finding that our techniques yield a consistent picture of the reasoning trace’s structure. We provide an open-source tool (thought-anchors.com) for visualizing the outputs of our methods on further problems. The convergence across our methods shows the potential of sentence-level analysis for a deeper understanding of reasoning models.
@@ 1 Introduction
Training large language models to reason with chain-of-thought (Reynolds & McDonell, 2021; Nye et al., 2021; Wei et al., 2023) has led to significant advances in capabilities (OpenAI, 2024). The resulting reasoning traces are regularly used in safety research (Baker et al., 2025; Shah et al., 2025), but there has been little work adapting interpretability methods to this new paradigm (though see (Venhoff et al., 2025; Goodfire, 2025)). Traditional mechanistic interpretability (Olah et al., 2020; Olah, 2022) methods often focus on a single forward pass of the model: understanding layer-by-layer activations and how these translate into a final output (Wang et al., 2022; Heimersheim & Janiak, 2023; Hanna et al., 2023). However, this framework is too fine-grained for autoregressive reasoning models, which consume their own output tokens.
@@
A core step in many interpretability strategies is to decompose the model into smaller parts that can be analyzed independently (Lindsey et al., 2025). A natural decomposition for chain-of-thought is into individual sentences and how they depend on each other. Interpretations of neural network behavior operate at varying levels of abstraction (Geiger et al., 2021; 2025), and sentence-level explanations strike an intermediate abstraction depth. Compared to tokens, sentences are more coherent and often coincide with reasoning steps extracted by an LLM (Venhoff et al., 2025; Arcuschin et al., 2025). Compared to paragraphs, sentences are less likely to conflate reasoning steps and may serve as an effective target for linking different steps.
Figure 1: Summary of our methods for principled attribution to important sentences in reasoning traces. A. An example reasoning trace with sentences labeled per our taxonomy. B. Our proposed methods are: black-box resampling, receiver heads, and attention suppression. C. A directed acyclic graph among sentences prepared by one of our techniques, made available open source.
@@
Prior work has established that CoT contains reasoning steps performing distinct functions. Backtracking sentences (e.g., “Wait…”) cause the model to revisit earlier conclusions, which boosts answer accuracy (Muennighoff et al., 2025). Other research has distinguished sentences based on whether they retrieve new information or execute deduction with existing information (Venhoff et al., 2025). Hence, reasoning may follow an overarching structure, where computational goals are generated, revised, and pursued. Yet, approaches for mapping this high-level structure are limited.
We propose that reasoning traces can be understood through thought anchors: critical reasoning steps that guide the trajectory of reasoning. We provide evidence for this type of anchoring based on black-box evidence from resampling and white-box evidence based on attention patterns. By measuring the causal dependencies between sentences via a masking approach, we further show how a CoTs wider computational structure can be interpreted. These measures go beyond just reading a CoT’s text, providing a principled foundation for interpretability that sidesteps disputes about the “faithfulness” of CoT text (Turpin et al., 2023; Korbak et al., 2025).
@@
Section˜2 and Section˜3 introduce a black-box method for measuring the counterfactual impact of a sentence on the model’s final answer and future sentences. We repeatedly resample reasoning traces from the start of each sentence. Based on resampling data, we can quantify the counterfactual impact of each sentence on the likelihood of any final answer. Additionally, we can distinguish planning sentences that initiate computations leading to some answer from sentences performing computations necessary for the answer but which are predetermined. Section˜4, adds a white-box method for evaluating importance based on the sentences most attended. Our analyses reveal “receiver” heads that narrow attention toward particular past “broadcasting” sentences. This provides a mechanistic measure of importance, whose findings converge with our resampling technique.
@@
Section˜5 and Section˜6 present a method mapping the wider structure of the reasoning the causal dependencies between pairs of CoT sentences. For each sentence in a trace, we intervene by masking all attention to it from subsequent tokens or simply removing the sentence entirely. We then measure the effect on subsequent token logits (KL divergence) compared to those generated without masking. Averaging token effects by sentence, this strategy measures each sentence’s direct causal effect on each subsequent sentence.
@@
2 Quantifying sentence importance
2.1 Model and dataset
Our analyses of sentence importance are based on the DeepSeek R1-Distill Qwen-14B model (48 layers) (DeepSeek, 2025). We used a temperature of $0.6$ and a top-p value of $0.95$. We focus on the MATH dataset (Hendrycks et al., 2021). Our analysis hinges on variability in final responses, so we target 20 challenging but doable questions that are correctly solved 25-75% of the time, identified by testing on 1,000 problems 10 times each. For each selected problem, we generated one correct and one incorrect reasoning trace, producing 40 responses. The average response is 144.2 sentences (95% CI: [116.7, 171.8]) and 4208 tokens (95% CI: [3479, 4937]). We focus only on sentences before the model has converged on an answer (i.e., after which it gives the same response in >98% of resamples). In Appendix˜B, we provide results from applying our techniques to the R1-Distill-Llama-8B model.
2.2 Forced answer importance
[…]
2.3 Importance via resampling
A limitation of the forced-answer approach is that a sentence $S$ may be necessary for some final answer but is consistently produced by the LLM late in the reasoning trace (e.g., a reliable arithmetic statement). This means that forced answer accuracy will be low for all sentences before $S$, precluding earlier step importance from being assessed.
Our approach evaluates importance by examining how a sentence may guide downstream sentences. Consider a rollout consisting of sentences $S_{1},S_{2},\dots,S_{i},\dots,S_{M}$ and a final answer $A$. We can use resampling to capture the extent sentence $S$ influences $A$. Specifically, for a given sentence $S_{i}$, we generate a distribution over final answers by generating 100 rollouts both without sentence $S_{i}$ (the base condition, with rollouts of the form $S_{1},S_{2},\dots,S_{i-1},T_{i},\dots,T_{N},A^{\prime}_{S_{i}}$), and another distribution with sentence $S_{i}$ (the intervention condition, with rollouts of the form $S_{1},S_{2},\dots,S_{i-1},S_{i},\dots,S_{M},A_{S_{i}}$). To assess the utility of resampling, we first conducted a brief case study.
2.4 Case study
We first investigate the efficacy of our sentence importance technique by applying it to one problem: “When the base-16 number $66666_{16}$ is written in base 2, how many base-2 digits (bits) does it have?” (MATH Problem 4682; see Section˜C.1 for the CoT transcript). The resampling data shows that from sentences 6-12, expected accuracy steadily declines, but sentence 13 causes accuracy to drastically increase (indicated by the navy and red circles in Figure 2A).
Figure 2: Accuracy over 100 rollouts at each (left) token, (middle) sentence, or (right) paragraph. Navy and red circles border the most importance sentence (Sentence 13) and are plotted in each graph as a reference. For the token graph, resampling was only done on the first 1,000 tokens of the CoT.
The large accuracy fluctuation motivates inspection of this part of the CoT. The model initially considers that $66666_{16}$ contains five base-16 digits, and any base-16 digit can be represented with four base-2 digits. Thus, the model considers the answer: 20 bits. However, this overlooks that $6_{16}$ is $110_{2}$ rather than $0110_{2}$ due to the leading zero. Interestingly, Sentence 12 mentions “checking if there’s any leading zero that might affect the bit count,” yet Sentence 12 lowers the expected accuracy. The uplift comes from Sentence 13, where the model decides to “calculate the value of $66666_{16}$ in decimal” (see resample alternatives in Section˜C.2). Downstream reasoning computes the decimal value of $66666_{16}$ and converts it to binary to arrive at the correct answer: 19 bits. The pivotal role of Sentence 13 is instead missed if one examines forced-accuracy importance (Figure 2A). This case study provides initial evidence that our resampling-based strategy identifies key moments in a CoT, where impactful plans are set or modified.
Further analyses show the efficacy of specifically examining sentences. The sentence-level resampling data mirrors the patterns seen resampling tokens but at a fraction of the cost (Figure 2B). Resampling paragraphs leads to meaningfully less resolution (Figure 2C). Although future work may benefit from defining reasoning steps at adaptive scales, the current results suggest that considering sentences provides high resolution while being cheaper than tokens.
3 Consistent patterns in sentence importance
3.1 Sentence taxonomy
To more systematically test whether reasoning is characterized by key sentences with outsized impacts, we organized sentences into different categories and measured their causal impacts. We adopted the framework by Venhoff et al. (2025), which describes distinct reasoning functions within a reasoning trace. We define eight categories (see examples and frequencies in Appendix˜D):
@@ 3.1 Sentence taxonomy
Each sentence in the analyzed response is assigned to one of these categories using an LLM-based auto-labeling approach (detailed in Appendix˜E). Categories that rarely appear are omitted from the figures below. Residual-stream probes accurately distinguish categories (see Appendix˜F).
@@ 3.2 Counterfactual importance
We additionally formalize our approach to quantifying importance in a manner that can be applied to any problem, including ones with any number of possible outcomes. We present two measures:
@@
    Resampling importance. We can compute the KL Divergence between the final answer distributions in the two conditions, i.e., $\text{importance}_{\text{r}}:=D_{\text{KL}}[p(A^{\prime}_{S_{i}})||p(A_{S_{i}})]$, providing a scalar measure of how much sentence $S_{i}$ changes the answer. We call this resampling importance. We include $\epsilon=10^{-9}$ to avoid division by zero, but the below conclusions remain consistent if instead performed using additive smoothing ($\alpha=0.5$ or $1.0$).
@@
    Counterfactual importance. The problem with resampling importance is that if $T_{i}$ is identical or similar to $S_{i}$ then we do not get much information about whether $S_{i}$ is important or not. Therefore, we write $S\not\approx T$ if two sentences $S$ and $T$ are dissimilar, defined as having embeddings with a cosine similarity less than the median value across all sentence pairs in our dataset (see Appendix˜G for details). Therefore, we can define counterfactual importance by conditioning on $T_{i}\not\approx S_{i}$; i.e., $\text{importance}:=D_{\text{KL}}[p(A^{\prime}_{S_{i}}|T_{i}\not\approx S_{i})||p(A_{S_{i}})]$.
@@
Because we resample all steps after a given sentence $S_{i}$, we avoid the aforementioned limitation of forced-answering. We also provide empirical evidence that the principled counterfactual importance definition in 1-3 above is useful, by comparing it to the resampling importance in Appendix˜I. As a comparison, we also evaluate forced answer importance based on this KL divergence strategy.
3.3 Results
@@
Plan generation and uncertainty management (e.g., backtracking) sentences consistently show higher counterfactual importance than other categories like fact retrieval or active computation (see Figure 3B). This supports the view that high-level organizational sentences anchor, organize, and steer the reasoning trajectory. These findings deviate from the analysis of forced answer importance, which instead implicates active computation as producing the greatest distributional shifts (Figure 3A). The forced-answer approach entirely neglects the importance of planning that influences other sentences, which we argue is more meaningful for understanding the trajectory of a reasoning trace.
@@
Figure 3: The mean of each sentence category for (A) forced-answer importance and (B) counterfactual importance, per the resampling method, plotted against the sentence category’s mean position in the reasoning trace. Only the 5 most common sentence types are shown (see Appendix˜H).
4 The mechanistic roots of importance
@@
We hypothesize that important sentences may receive heightened attention from downstream sentences. Although attention weights do not necessarily imply causal links (see Section 5), heightened attention is a plausible mechanism by which important sentences influence subsequent sentences. We conjecture further that a high focus on important sentences may be driven by specific attention heads, and by tracking such heads, we may pinpoint key sentences.
@@
We assessed the degree different heads narrow attention toward particular sentences. First, for each reasoning trace, we averaged each attention head’s token-token attention weight matrix to form a sentence-sentence matrix, where each element is the mean across all pairs of tokens between two sentences. Based on each attention matrix, we computed the mean of its columns below the diagonal to measure the extent each sentence receives attention from all downstream sentences; averaged only among pairs at least four sentences apart to focus on distant connections. This generates a distribution for each head (e.g., Figure 4A), and the extent each head narrows attention toward specific sentences can be quantified as its distribution’s kurtosis (computed for each reasoning trace, then averaged across traces). Plotting each head’s kurtosis reveals that some attention heads strongly narrow attention toward specific, possibly important, sentences in the reasoning trace (Figure 4B).